### v1

# Birack Bracket Quivers and Framed Links

Abstract
We introduce birack brackets, skein invariants of birack-colored framed classical and virtual knots and links with values in a commutative unital ring. The multiset of birack bracket values over the homset from a framed link’s fundamental birack then forms an invariant of framed links. We then categorify this multiset to define a quiver-valued invariant of framed knots and links. From this quiver we define new polynomial invariants of framed knots and links.
Keywords: Biracks, birack brackets, quantum enhancements of counting invariants, framed knots and links
2020 MSC: 57K12
1 Introduction
Biracks are algebraic structures whose axioms encode the framed version of the Reidemeister moves, first introduced in [9] and later studied in works such as [12, 15] etc. Biracks satisfying a stronger axiom coming from the unframed Reidemeister I move known as biquandles, notable for their utility in defining invariants of unframed knots and links, have been studied in many works such as [6, 8, 13] and more.
In particular, the birack homset invariant is the set of birack homomorphisms from the fundamental birack of an framed oriented classical or virtual knot or link $L$ to a (usually finite) birack $X$. The elements of this homset invariant can be represented as birack colorings of a diagram $D$ of $L$, i.e., assignments of elements of $X$ to the semiarcs in a diagram of $L$ satisfying a certain coloring condition at each crossing.
Invariants of birack-colored diagrams provide enhancements of the birack counting invariant, generally stronger invariants of framed oriented knots and links which determine the birack counting invariant. In some cases these can be summed over a finite set of framings of a knot or link to obtain invariants of unframed oriented knots and links; see [2, 4, 15] for examples.
Biquandle brackets are skein invariants of biquandle-colored diagrams with skein coefficients depending on the colors at a crossing, first introduced in [13] and later studied in works such as [10, 11, 14] etc. The biquandle bracket invariant was categorified to the case of biquandle bracket quivers in [7] by applying a construction from [5].
In this paper we generalize the biquandle bracket construction to the case of framed oriented classical or virtual knots and links, obtaining birack brackets and their categorifications, birack bracket quivers. The paper is organized as follows. In Section 2 we review the basics of biracks and the birack homset for framed links. In Section 3 we introduce the birack bracket construction along with a resulting polynomial invariant of framed links. In Section 4 we categorify the birack bracket invariant to obtain birack bracket quivers and use these to obtain further polynomial invariants of framed knots and links. We collect some examples and computations, and we conclude in Section 5 with some questions for future research.
This paper, including all text, illustrations, and python code for computations, was written strictly by the authors without the use of generative AI in any form.
2 Birack Review
As mentioned in the Introduction, to define knot invariants various mathematical structures have been implemented to encode the crossing information of a knot, taking the forms of groups, tensors, and more abstract algebraic structures. Reidemeister moves characterize the equivalence relation of ambient isotopy on diagrams; the corresponding algebraic structures in result in different knot invariants. More specifically, quandles are represent the algebraic structure of Reidemeister moves on arcs in classical knots; biquandles represent Reidemeister moves on semi-arcs in classical and virtual knots, and racks represent the algebraic structure of arcs in framed knots and links through replacing Reidemeister I move with the framed version of the move. In this paper, we consider the case of biracks, algebraic structures generated by semiarcs in a link diagram with axioms corresponding to blackboard framed isotopy.
Recall that two blackboard framed oriented knot or link diagram represent isotopic framed knots or links if and only if they are related by a sequence of the blackboard framed Reidemeister moves, of which a generating set are:
Next we recall a definition from [12], reformulated in our current notation.
Definition 1.
A framed birack is a set $X$ with two binary operations $\underline{\triangleright}\,$$\overline{\triangleright}\,$ satisfying the axioms
  • (i)
    For all $x\in X$ we have
    $(x\,\underline{\triangleright}\,x)\,\overline{\triangleright}\,(x\,\overline{\triangleright}\,x)=(x\,\overline{\triangleright}\,x)\,\underline{\triangleright}\,(x\,\underline{\triangleright}\,x),$
  • (ii)
    The maps $\alpha_{x},\beta_{x}:X\to X$ and $S:X\times X\to X\times X$ defined by $\alpha_{x}(y)=y\,\overline{\triangleright}\,x$, $\beta_{y}(x)=y\,\underline{\triangleright}\,x$ and $S(x,y)=(y\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,y)$ are invertible for all $x\in X$ and
  • (iii)
    The exchange laws
[…]
    are satisfied for all $x,y,z\in X$.
A birack in which we also have $x\,\underline{\triangleright}\,x=x\,\overline{\triangleright}\,x$ for all $x\in X$ is a biquandle. A biquandle in which $x\,\overline{\triangleright}\,y=x$ for all $x,y\in X$ is a quandle.
Remark 1.
Biracks are sometimes defined without requiring our axiom (i); we will refer to these as regular biracks. Essentially, regular biracks can be used to define invariants of framed isotopy, e.g. braid invariants, while framed biracks provide invariants of framed knots and links. Unless otherwise specified, we will mean “framed birack” whenever we say “birack”.
As observed in [12], for any finite birack $X$ there is a unique bijection $\pi:X\to X$ called the kink map satisfying
$x\,\underline{\triangleright}\,\pi(x)=\pi(x)\,\overline{\triangleright}\,x$
for all $x\in X$. A biquandle is a birack whose kink map is the identity.
Example 1.
One infinite family of examples of biracks is the set of $(t,s,r)$-biracks. Let $X$ be a $\widetilde{\Lambda}-$module where $\widetilde{\Lambda}=\mathbb{Z}[t^{\pm 1},s,r^{\pm 1}]/(s^{2}-s(r-t))$. Then $X$ is a birack under the operations
$x\,\overline{\triangleright}\,y=rx,\quad x\,\underline{\triangleright}\,y=tx+sy.$
To see this we can verify the axioms:
Checking axiom (i), we have
$(x\,\underline{\triangleright}\,x)\,\overline{\triangleright}\,(x\,\overline{\triangleright}\,x)$ $=$ $r(tx+sx)$
$=$ $r(t+s)x$
$=$ $(tr+st+s(r-t))x$
$=$ $(tr+st+s^{2})x$
$=$ $t(rx)+s(tx+sx)$
$=$ $(x\,\overline{\triangleright}\,x)\,\underline{\triangleright}\,(x\,\underline{\triangleright}\,x)$
as required.
Axiom (ii) says that $S(x,y)=(y\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,y)=(ry,tx+sy)$ is invertible for a birack; namely, there exists $S^{-1}$ such that $S^{-1}(y\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,y)=(x,y)$, or $S^{-1}(ry,tx+sy)=(x,y)$. Then we note that $S^{-1}$ is given by $S^{-1}(x,y)=(t^{-1}(y-sr^{-1}x),r^{-1}x)$, where $s$ is not necessarily invertible in $\widetilde{\Lambda}$.
Finally, checking the the exchange laws of axiom (iii):
$(x\,\underline{\triangleright}\,y)\,\underline{\triangleright}\,(z\,\underline{\triangleright}\,y)$ $=$ $(tx+sy)\,\underline{\triangleright}\,(tz+sy)$
$=$ $t^{2}x+sty+stz+s^{2}y$
$=$ $=t^{2}x+tsz+rsy$
$=$ $(x\,\underline{\triangleright}\,z)\,\underline{\triangleright}\,(y\,\overline{\triangleright}\,z)$
since $s^{2}+st=rs$,
$(x\,\underline{\triangleright}\,y)\,\overline{\triangleright}\,(z\,\underline{\triangleright}\,y)$ $=$ $(tx+sy)\,\overline{\triangleright}\,(tz+sy)$
$=$ $r(tx+sy)$
$=$ $rx\,\underline{\triangleright}\,ry$
$=$ $(x\,\overline{\triangleright}\,z)\,\underline{\triangleright}\,(y\,\overline{\triangleright}\,z)$
and
$(x\,\overline{\triangleright}\,y)\,\overline{\triangleright}\,(z\,\overline{\triangleright}\,y)$ $=$ $rx\,\overline{\triangleright}\,rz$
$=$ $r^{2}x$
$=$ $rx\,\overline{\triangleright}\,(ty+sz)$
$=$ $(x\,\overline{\triangleright}\,z)\,\overline{\triangleright}\,(y\,\underline{\triangleright}\,z)$
We note that the kink map in an Alexander birack is given by $\pi(x)=(r-s)t^{-1}x$; this is bijective since (rs)(r-s) is invertible in $\widetilde{\Lambda}$ with inverse $t^{-1}r^{-1}(s+t)$. Hence, any $\widetilde{\Lambda}-$module is an example of a birack with these operations.
Example 2.
One concrete example of a $(t,s,r)$-birack is $(t=1,s=2)$ on $\mathbb{Z}_{4}$. By the condition $s^{2}=s(r-t)$, $r$ is determined to be either 3 or 1; let us choose $r=1$. The operation tables for $\underline{\triangleright}\,$$\overline{\triangleright}\,$ are then given below:
$\begin{array}[]{c|cccc}\,\underline{\triangleright}&0&1&2&3\\ \hline\cr 0&0&2&0&2\\ 1&1&3&1&3\\ 2&2&0&2&0\\ 3&3&1&2&1\\ \end{array}\quad\begin{array}[]{c|cccc}\,\overline{\triangleright}&0&1&2&3\\ \hline\cr 0&0&0&0&0\\ 1&1&1&1&1\\ 2&2&2&2&2\\ 3&3&3&3&3\end{array}.$
The kink map in this case is $\pi(x)=3x$.
Definition 2.
Let $X$ and $Y$ be biracks with operations $\underline{\triangleright}\,_{X},\,\overline{\triangleright}\,_{X},\,\underline{\triangleright}\,_{Y},\,\overline{\triangleright}\,_{Y}$. A birack homomorphism is a function $f:X\to Y$ such that for all $x,y\in X$ we have
$f(x\,\underline{\triangleright}\,_{X}\,y)=f(x)\,\underline{\triangleright}\,_{Y}f(y)\quad\mathrm{and}\quad f(x\,\overline{\triangleright}\,_{X}\,y)=f(x)\,\overline{\triangleright}\,_{Y}f(y).$
Definition 3.
A birack coloring of a framed oriented knot or link diagram is an assignment of elements of our birack $X$ to the semiarcs of our diagram such that at every crossing we have
[…]
The birack axioms are chosen so that for every birack coloring of a framed oriented link diagram before a framed Reidemeister move, there is a unique coloring of the diagram after the move that agrees with the original coloring outside the neighborhood of the move.
Definition 4.
Let $L$ be a framed oriented knot or link diagram. The fundamental birack of $L$, denoted $\mathcal{B}(L)$, has a presentation with generators corresponding to the semiarcs in $L$ and relations between the generators at crossings as depicted in Definition 3. The elements of $\mathcal{B}(L)$ are then equivalence classes of birack words in these generators (including symbols such as $\alpha_{x}^{-1}(y)$ etc. as required by the axioms) modulo the equivalence relation generated by the crossing relations and the biquandle axioms.
A birack coloring of a framed knot diagram $L$ by a finite birack $X$ determines a birack homomorphism $f:\mathcal{B}(L)\to X$ by assigning an image to each generator consistently with the birack axioms. That is, a birack homomorphism $f$ assigns to each semi-arc in $\mathcal{B}(L)$ an image element in $X$ so that the crossing information of semi-arcs in $\mathcal{B}(L)$ satisfies the algebraic relations of the birack $X$. One framed knot birack homomorphism represents one coloring of the framed link $L$. The set of all framed knot birack homomorphisms from the fundamental birack of $L$ to a birack $X$, known as the birack homset, can then be identified with the set of all possible colorings of $K$, denoted by $\mathrm{Hom}\mathcal{(}{B}(L),X)$.
Definition 5.
The birack counting invariant of a framed oriented knot or link $L$ is defined as the cardinality of the homset,
$\Phi_{X}^{\mathbb{Z}}=|\mathrm{Hom}(\mathcal{B}(L),X)|.$
Example 3.
The fundamental birack of the framed knot $3_{1}$ can be computed by labeling semi-arcs of the knot as generators and crossing information as relations:
$\mathcal{B}(3_{1})=\bigl\langle\,x,y,z,u,v,w\;\bigm|\;x\,\underline{\triangleright}\,y=u,\;y\,\overline{\triangleright}\,x=w,\;y\,\underline{\triangleright}\,z=v,\;z\,\overline{\triangleright}\,y=u,\;z\,\underline{\triangleright}\,x=w,\;x\,\overline{\triangleright}\,z=v\bigr\rangle.$
Then let $X$ be the birack given by the operation tables
[…]
All possible homomorphisms from the fundamental birack $\mathcal{B}(3_{1})$ to the birack $X$ form the homset $\mathrm{Hom}(\mathcal{B}(3_{1}),X)$. There are three $X$-colorings:
Therefore, the birack counting invariant $|\mathrm{Hom}(\mathcal{B}(3_{1}),X)|$ is 3.
Remark 2.
In [12] the birack counting invariant for framed oriented knots and links was used to define an invariant of unframed oriented knots and links by collecting the counting invariant values of a complete period of framings modulo the birack rank, i.e., the smallest integer $n\geq 1$ such that $\pi^{n}=\mathrm{Id}$.
3 Birack Brackets
We now state our main definition.
Definition 6.
Let $X$ be a birack and $R$ a commutative unital ring. A birack bracket is defined by two maps $A,B:X\times X\to R^{\times}$ denoted $A(x,y)=A_{x,y}$ and $B(x,y)=B_{x,y}$ such that
  • (i)
    For all $x\in X$ we have
    $A_{x,x}^{2}B_{x,x}^{-1}=A_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}^{2}B_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}^{-1},$
  • (ii)
    For all $x,y\in X$,
    $\delta=-A_{x,y}B_{x,y}^{-1}-A_{x,y}^{-1}B_{x,y}\neq 0$
    and
  • (iii)
    For all $x,y,z\in X$,
[…]
where $A(x,y)$ and $B(x,y)$ are denoted $A_{x,y}$ and $B_{x,y}$ and $A_{x,x}^{2}B_{x,x}^{-1}$ is denoted by $w_{x}$. A birack in which $w_{x}=w_{y}$ for all $x,y\in X$ is called homogeneous.
The maps $A,B:X\times X\to R^{\times}$ give the smoothing coefficients using the birack skein relations:
[…]
The birack bracket axioms are then the conditions required for the state-sum value to be unchanged by framed Reidemeister moves. More precisely, simultaneously performing all smoothings results in a sum of products of skein coefficients times diagrams without crossings, i.e. disjoint closed curves, known as Kauffman states. The value of a state is $\delta^{k}$ where $k$ is the number of disjoint curves in the state times the associated product of skein coefficients. Indeed, we have:
Proposition 1.
Let $L$ be an oriented link diagram, $X$ a finite birack and $\beta$ a birack bracket with coefficients in a commutative unital ring $R$. Then for each coloring $f\in\mathrm{Hom}(\mathcal{BR}(L),X)$, the state-sum value
$\beta(f)=\sum_{\mathrm{Kauffman\ states}}\left(\prod_{\mathrm{smoothings}}C_{xy}^{\pm 1}\delta^{k}\right)$
defined by summing over all Kauffman states the product of the state’s smoothing coefficients $C_{xy}^{\pm 1}$ times $\delta$ to the power of the number of components in the state is invariant under $X$-colored framed Reidemeister moves.
Proof.
Invariance under the direct and reverse type II moves and the all-positive type III move is the same as in [13]; let us consider the framed type I move.
Comparing the two sides of the move we obtain the requirement that
$A_{xx}\delta+B_{xx}=A_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}\delta+B_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}$
which says
$A_{xx}(-A_{xx}B_{xx}^{-1}-A_{xx}^{-1}B_{xx})+B_{xx}=A_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}(-A_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}B_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}^{-1}-A_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}^{-1}B_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x})+B_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}$
which reduces to
$A_{xx}^{2}B_{xx}^{-1}=A_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}^{2}B_{x\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,x}^{-1}.$
We can specify a birack bracket for a finite biquandle $X$ with values in $R$ with a block matrix $[A|B]$ whose entries are the smoothing coefficients $A_{i,j},B_{i,j}$.
Example 4.
The birack $X$ specified by the operation tables
[…]
has birack brackets over $\mathbb{Z}_{5}$ including
$\left[\begin{array}[]{rrr|rrr}1&1&3&4&4&2\\ 3&2&4&2&3&1\\ 1&2&3&4&3&2\end{array}\right]$
and over $\mathbb{C}$ including
$\left[\begin{array}[]{rrr|rrr}1&-i&-i&-1&i&i\\ -i&1&-1&i&-1&1\\ 1&1&-i&-1&-1&i\end{array}\right].$
Corollary 2.
Let $L$ be an oriented link diagram, $X$ a finite birack and $\beta$ a birack bracket with coefficients in a commutative unital ring $R$. Then the multiset of $\beta$ values over the birack homset, denoted $\Phi_{X}^{\beta,M}(L)$, is an invariant of oriented framed knots and links.
Definition 7.
We can convert the birack bracket multiset invariant into a polynomial for ease of comparison. We define the birack bracket polynomial associated to a birack $X$ and birack bracket $\beta$ as
$\Phi_{X}^{\beta}(L)=\sum_{x\in\Phi_{X}^{\beta,M}(L)}u^{x}.$
Example 5.
Consider the birack over the set $X=\{1,2,3\}$ with operation tables given by
$\begin{array}[]{r|rrr}\,\underline{\triangleright}&1&2&3\\ \hline\cr 1&2&2&2\\ 2&3&3&3\\ 3&1&1&1\\ \end{array}\quad\begin{array}[]{r|rrr}\,\overline{\triangleright}&1&2&3\\ \hline\cr 1&3&3&3\\ 2&1&1&1\\ 3&2&2&2\end{array}.$
One birack bracket with coefficients in $\mathbb{Z}_{5}$ as computed with our python code is given by
[…]
with cycle value
$\delta=-A_{x,y}B_{x,y}^{-1}--A_{x,y}^{-1}B_{x,y}=-(1)(1)-(1)(1)=-2=3$
which is the same for all values of $x$ and $y$.
Applying the birack bracket skein relation to obtain the state sum for the framed trefoil knot with writhe 3, we have:
yielding
$\beta(f)$ $=$ $A_{xy}A_{zx}A_{yz}\delta^{2}+A_{xy}A_{zx}B_{yz}\delta+A_{xy}B_{zx}A_{yz}\delta+A_{xy}B_{zx}B_{yz}\delta^{2}$
$+B_{xy}A_{zx}A_{yz}\delta+B_{xy}A_{zx}B_{yz}\delta^{2}+B_{xy}B_{zx}A_{yz}\delta^{2}+B_{xy}B_{zx}B_{yz}\delta^{3}$
Then substituting in the values of $A_{xy},B_{xy}$ etc. for each of the three colorings, we have
$\begin{array}[]{ccc|c}x&y&z&\phi\\ \hline\cr 2&3&1&(3)(2)(3)(3^{2})\times 4+(3)(2)(3)(3)\times 3+(3)(2)(3)(3^{2})=3+2+1=1\\ 3&1&2&(2)(3)(3)(3^{2})\times 4+(2)(3)(3)(3)\times 3+(2)(3)(3)(3^{2})=3+2+1=1\\ 1&2&3&(3)(3)(2)(3^{2})\times 4+(3)(3)(2)(3)\times 3+(3)(3)(2)(3^{2})=3+2+1=1\end{array}$
The resulting birack bracket multiset of the trefoil knot is $\Phi_{X}^{\beta,M}(3_{1})$ is $\{1,1,1\}$, which in polynomial form is $3u$.
Remark 3.
The framed trefoil with writhe 4 has no colorings by the birack in Example 5 and hence $\Phi_{X}^{\beta,M}(L)=\emptyset$ which in polynomial form is 0. In particular the birack bracket invariant can distinguish different framings of the same framed knot.
Example 6.
Let $X$ be the birack given by the operation tables
$\begin{array}[]{r|rrr}\,\underline{\triangleright}&1&2&3\\ \hline\cr 1&2&2&1\\ 2&1&1&2\\ 3&3&3&3\\ \end{array}\quad\begin{array}[]{r|rrr}\,\underline{\triangleright}&1&2&3\\ \hline\cr 1&1&1&2\\ 2&2&2&1\\ 3&3&3&3\\ \end{array}.$
We compute via python that the coefficient matrix
$\left[\begin{array}[]{rrr|rrr}1&1&-1&-1&-1&1\\ -1&-1&-1&1&1&1\\ i&-i&1&-i&i&1\end{array}\right]$
defines a birack bracket on $X$ with coefficients in $\mathbb{C}$. Then the framed Hopf link $L$
has birack bracket polynomial
$\Phi_{X}^{\beta}(L)=4u^{4i}+5u^{4}.$
4 Birack Bracket Quivers
We begin this Section with a definition generalizing a definition in [7].
Definition 8.
Let $X$ be a finite birack, $S\subseteq\mathrm{Hom}(X,X)$ a set of birack endomorphisms and $L$ an oriented framed link. Then the birack coloring quiver of $L$ with respect to $X$ and $S$, denoted $\mathcal{BRQ}_{X,S}(L)$, is the directed graph with a vertex for each birack coloring $v\in\mathrm{Hom}(\mathcal{BR}(L),X)$ and a directed edge from vertex $v_{1}$ to vertex $v_{2}$ whenever $v_{2}=v_{1}\sigma$ for each endomorphism $\sigma\in S$.
This definition is motivated by the observation that if $\sigma:X\to X$ is a birack endomorphism and $v_{1}$ is a valid birack coloring of a diagram of $L$, then applying $\sigma$ to each color in $v_{1}$ yields another valid coloring, $v_{2}=v_{1}\sigma$.
As in previous cases such as [3, 5, 7] etc., we observe that since the ingredients for $\mathcal{Q}_{X,S}(L)$ are all invariant under framed Reidemeister moves, so is $\mathcal{Q}_{X,s}(L)$. That is, we have:
Theorem 3.
For any finite biquandle $X$ and $S\subseteq\mathrm{Hom}(X,X)$, the quiver $\mathcal{Q}_{X,S}(L)$ is an invariant of framed isotopy.
Next, following earlier work such as [7], we come to our primary new definition.
Definition 9.
Let $X$ be a finite birack, $S\subseteq\mathrm{Hom}(X,X)$ a set of birack endomorphisms, $\beta$ a birack bracket and $L$ an oriented framed link. The birack bracket quiver of $L$ with respect to $X,S$ and $\beta$, denoted $\mathcal{BRQ}_{X,\beta}^{S}(L)$, is the birack coloring quiver of $L$ with respect to $X$ and $S$ with vertices $v$ weighted with the birack bracket values $\beta(v)$ associated to the colorings $v$.
A finite quiver is a small category with vertices as objects and directed paths as morphisms; hence, Definition 9 establishes a family of categorifications of the birack bracket multiset and polynomial invariants from the previous section.
The isomorphism class of $\mathcal{BRQ}_{X,S}^{\beta}$ is an invariant of framed oriented knot and links and can thus be used to distinguish framed oriented knots and links. Since these quivers become complex quite quickly, it is also useful to extract some simpler invariants via decategorification.
Definition 10.
Given a finite birack $X$, a framed knot $K$ and a birack bracket $\beta$ with coefficients in a commutative unital ring $R$ and $S\subset\mathrm{Hom}(X,X)$ a set of endomorphisms of $X$. Then the birack bracket quiver $\mathcal{BRQ}_{X,S}^{\beta}(K)$ has vertex set $V$ and edge set $E$. Let $P$ the set of maximal non-repeating paths in the quiver. We then decategorify this birack bracket quiver $\mathcal{BRQ}_{X,S}^{\beta}(K)$ to obtain the birack bracket in-degree polynomial
$\Phi_{X,\beta}^{S,\mathrm{deg}_{+}}(K)=\sum_{v\in V}u^{\beta(v)}v^{deg+(v)},$
the birack bracket two-variable polynomial
$\Phi_{X,\beta}^{S,2}(K)=\sum_{e\in E}s^{\beta(s(e))}t^{\beta(t(e))}$
where $s(e)$ and $t(e)$ represent the source and tail of edge $e\in E$, and the birack bracket maximal path polynomial
$\Phi_{X,\beta}^{S,MP}(K)=\sum_{p\in P}x^{\sum_{v\in p}\beta(v)}y^{|p|}$
where $v\in p$ means $v$ is a vertex in the path $p$ and $|p|$ is the length of the path $p$.
Example 7.
Recall the three different colorings of the trefoil knot $3_{1}$ by the constant action birack structure on $X=\{1,2,3\}$ in 3. Let the endomorphism $S=\{\phi\}\subset\mathrm{Hom}(X,X)$ where $\phi(1)=2,\phi(2)=3$ and $\phi(3)=1$. We can specify this endomoprhism with a list of output values, i.e. $\sigma=[2,3,1]$. Then we have birack coloring quiver
[…]
The birack coloring quiver then replaces each coloring with a vertex weighted by its birack bracket value; in this case we have
[…]
The framed knot invariant polynomials obtained from decategorification of this birack bracket quiver are then
$\Phi_{X,\beta}^{S,deg+}(3_{1})=3uv$
and
$\Phi_{X,\beta}^{S,2}(3_{1})=3st.$
Example 8.
The birack in Example 6 has endomorphism set $\mathrm{Hom}(X,X)=\{\sigma_{1},\sigma_{2},\sigma_{3}\}=\{[1,2,3],[3,3,3],[2,1,3]\}$, i.e., where $\sigma_{1}$ is the identity map, $\sigma_{2}$ is the constant map sending everything to 3, and $\sigma_{3}$ fixes 3 while swapping 1 and 2. Then the birack bracket quiver of the framed Hopf link in Example 6 with respect to the birack bracket in the same example and endomorphism set $S=\{\sigma_{2},\sigma_{3}\}$ is
[…]
Example 9.
Let $X$ be the birack defined by the operation tables
[…]
Then one birack bracket found with our python code over $\mathbb{Z}_{5}$ is given by
$\left[\begin{array}[]{cccc|cccc}1&1&2&1&2&2&4&2\\ 2&2&2&1&4&4&4&2\\ 1&1&1&3&2&2&2&1\\ 2&2&4&2&4&4&3&4\end{array}\right].$
The set of endomorphisms of this birack $X$ is
$\mathrm{Hom}(X,X)=\{[4,2,3,1],[4,3,2,1],[1,2,3,4],[1,3,2,4]\}.$
With these, the birack bracket quiver invariants for various blackboard framed links from the table at [1] are:
$\begin{array}[]{r|l}\Phi_{X,\beta}^{S,2}(L)&L\\ \hline\cr 16u^{4}v^{4}&L2a1,L6a2,L6a3,L7a5,L7a6\\ 16u^{4}v^{4}+32u^{3}v^{4}+16uv^{4}&L4a1\\ 32u^{4}v^{4}&L5a1,L7a1,L7a3,L7a4,L7n2\\ 32u^{4}v^{4}+32u^{3}v^{4}&L6a1\\ 32u^{4}v^{4}+32u^{3}v^{4}+96u^{2}v^{4}+96uv^{4}&L6a4\\ 32u^{4}v^{4}+32u^{3}v^{4}+96u^{2}v^{4}&L6a5,L6n1\\ 16u^{4}v^{4}+16u^{3}v^{4}&L7a2\\ 32u^{3}v^{4}+32uv^{4}&L7a7\\ 16u^{4}v^{4}+16u^{2}v^{4}&L7n1\\ \end{array}.$
With the same birack and birack bracket, the birack bracket quiver invariants for various blackboard framed virtual knots from the table at [1] are:
$\begin{array}[]{r|l}\Phi_{X,\beta}^{S,2}(VK)&VK\\ \hline\cr 8u^{4}v^{4}+8u^{2}v^{4}&2.1,4.21,4.24,4.28,4.36\\ 8u^{2}v^{4}&3.1,3.3,3.3,3.4,3.5,3.6,3.7\\ 8u^{3}v^{4}+8u^{2}v^{4}&4.1,4.3,4.9,4.15,4.25,4.28,4.29,4.37\\ 16u^{2}v^{4}&4.2,4.6,4.8,4.12,4.13,4.14\\ 8u^{2}v^{4}+8uv^{4}&4.4,4.5,4.10,4.11,4.16,4.18,4.23,4.27,4.30,4.31,4.33,4.38,4.39\\ 16u^{2}v^{4}&4.2,4.17,4.19,4.20,4.22,4.26,4.32,4.34,4.35,4.40\end{array}$
Example 10.
Continuing with the birack and birack bracket from Example 9, the following example shows our new decategorification invariants distinguishing different framings of the same knot, in this case the Figure 8 knot $4_{1}$ with framings from 4-4 to $4$. We illustrate the cases of framing numbers 1-1, 0 and $1$:
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Then we compute invariant values in the table.
$\begin{array}[]{r|l}\Phi_{X,\beta}^{S,2}(4_{1})&4_{1}\mathrm{\ with\ framing\ number\ from\ -4\ to\ 4}\\ \hline\cr 16u^{2}v^{4}&4_{1}^{0}\\ 8u^{2}v^{4}&4_{1}^{1},4_{1}^{3},4_{1}^{-1},4_{1}^{-3}\\ 8u^{4}v^{4}+8u^{2}v^{4}&4_{1}^{2}\\ 8u^{3}v^{4}+8u^{2}v^{4}&4_{1}^{4},4_{1}^{-4}\\ 8u^{2}v^{4}+8uv^{4}&4_{1}^{-2}\end{array}$
where the superscript denotes framing number from 4-4 to $4$.
We conclude this section with another example.
Example 11.
Let $X$ be the birack given by the operation tables
$\begin{array}[]{r|rrr}\,\underline{\triangleright}&1&2&3\\ \hline\cr 1&2&2&2\\ 2&1&1&1\\ 3&3&3&3\end{array}\quad\begin{array}[]{r|rrr}\,\overline{\triangleright}&1&2&3\\ \hline\cr 1&1&1&2\\ 2&2&2&1\\ 3&3&3&3\end{array}.$
Then we compute via python that the block matrix
$\beta=\left[\begin{array}[]{rrr|rrr}1&1&i&-q&-q&iq\\ i&i&-1&-iq&-iq&q\\ i&1&1&-iq&-q&-q\end{array}\right]$
defines a birack bracket over $\mathbb{C}[q^{\pm 1}]$ and the maps
$\sigma_{1}=[3,3,3],\quad\sigma_{2}=[2,1,3]$
are endomorphisms of $X$. Then we compute that the framed virtual links
are distinguished by the birack maximal path polynomial with
$\Phi_{X,S}^{\beta,MP}(L_{1})=8x^{6q^{8}+6q^{2}+6+6q^{-2}}y^{5}+8y^{5}$
and
$\Phi_{X,S}^{\beta,MP}(L_{2})=4x^{(3-3i)q^{8}+(3-3i)q^{2}+3-3i+(3-3i)q^{-2}}y^{5}+4x^{(3+3i)q^{8}+(3+3i)q^{2}+3+3i+(3+3i)q^{-2}}y^{5}.$
We note that these two virtual links have the same Jones polynomial value of $q^{8}+q^{4}+1q^{-12}$, so this example shows that the birack bracket quiver (and hence its decategorifications) are not determined by the Jones polynomial.
5 Questions
It is important to note that the examples in this paper are, generally speaking, “toy” examples, using small-cardinality biracks and typically finite rings since these are easy to compute; they are not to be taken to represent the full power of this infinite family of invariants. In particular, larger biracks and larger finite or infinite coefficient rings should provide more powerful invariants.
In biquandle bracket theory, as in the Kauffman bracket polynomial, we can adjust the value of the bracket of any framing of a knot or link back to the value of the zero framing by multiplying by an appropriate power of $w$. In birack bracket theory this no longer works for the simple reason that different framings can have different numbers of colorings, so adjusting a given coloring to a given writhe may not make sense. This is reflected in the fact that the value of $w_{x}$ is not always invariant over the entire birack bracket. What do these values of $w_{x}$ imply about the algebraic structure of generic birack brackets? What global information about chosen biracks can a combination of $w_{x}$ values over corresponding birack brackets describe?
We are curious about the possibility of functors between framed knots/links and unframed knots/links. Over chosen finite rings on which the birack brackets are defined, one can add different framing numbers as shown in Example 10. The reverse process is to “smooth out” the framing information as a state sum of different framed knot/link states. The corresponding factors might be defined with respect to the $w$ values. We hope to resolve these issues in a future paper.
What information about a framed knot or link is contained in the geometric structure of the birack coloring quiver? We can observe that paths in the quiver always terminate in a cycle since the out-degree of every vertex is always positive; similarly, endomorphisms which are automorphisms tend to produce cycles in the quiver. What other characterizations can be made?
Finally, we ask what is the algebraic structure of the set of birack brackets? What operations can be used to combine brackets over a given birack to obtain valid birack brackets?

### v2

@@
 Reidemeister moves characterize the equivalence relation of ambient isotopy on diagrams; the corresponding algebraic structures in result in different knot invariants.
-More specifically, quandles are represent the algebraic structure of Reidemeister moves on arcs in classical knots; biquandles represent Reidemeister moves on semi-arcs in classical and virtual knots, and racks represent the algebraic structure of arcs in framed knots and links through replacing Reidemeister I move with the framed version of the move.
+More specifically, quandles represent the algebraic structure of Reidemeister moves on arcs in classical knots; biquandles represent Reidemeister moves on semi-arcs in classical and virtual knots, and racks represent the algebraic structure of arcs in framed knots and links through replacing Reidemeister I move with the framed version of the move.
 In this paper, we consider the case of biracks, algebraic structures generated by semiarcs in a link diagram with axioms corresponding to blackboard framed isotopy.
-Recall that two blackboard framed oriented knot or link diagram represent isotopic framed knots or links if and only if they are related by a sequence of the blackboard framed Reidemeister moves, of which a generating set are:
+Recall that two blackboard framed oriented knot or link diagram represent isotopic framed knots or links if and only if they are related by a sequence of the blackboard framed Reidemeister moves, of which a generating set is:
 Next we recall a definition from [12], reformulated in our current notation.
@@
   • (ii)
-    The maps $\alpha_{x},\beta_{x}:X\to X$ and $S:X\times X\to X\times X$ defined by $\alpha_{x}(y)=y\,\overline{\triangleright}\,x$, $\beta_{y}(x)=y\,\underline{\triangleright}\,x$ and $S(x,y)=(y\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,y)$ are invertible for all $x\in X$ and
+    The maps $\alpha_{x},\beta_{x}:X\to X$ and $S:X\times X\to X\times X$ defined by $\alpha_{x}(y)=y\,\overline{\triangleright}\,x$, $\beta_{y}(x)=x\,\underline{\triangleright}\,y$ and $S(x,y)=(y\,\overline{\triangleright}\,x,x\,\underline{\triangleright}\,y)$ are invertible for all $x\in X$ and
   • (iii)
@@
 Biracks are sometimes defined without requiring our axiom (i); we will refer to these as regular biracks.
-Essentially, regular biracks can be used to define invariants of framed isotopy, e.g.
+Essentially, regular biracks can be used to define invariants of regular isotopy, e.g.
 braid invariants, while framed biracks provide invariants of framed knots and links.
@@
 for all $x\in X$. A biquandle is a birack whose kink map is the identity.
+More precisely, the framed birack axioms are chosen so that for any valid birack coloring of an oriented framed link diagram on one side of a framed Reidemeister move, there is a unique valid coloring of the diagram on the other side of the move which agrees outside the neighborhood of the move.
+Then starting with a single strand, using Reidemeister II and III moves we can perform what’s sometimes called the Fenn-Rourke move (FR):
+[…]
+The bijectivity of colorings establishes the bijectivity of the kink map $\pi$; the framed Reidemeister I move condition expressed algebraically by axiom (i) then shows that the kink map (and hence its inverse) does not depend on which side of the strand the loop is on.
 Example 1.
@@
 […]
-The birack axioms are chosen so that for every birack coloring of a framed oriented link diagram before a framed Reidemeister move, there is a unique coloring of the diagram after the move that agrees with the original coloring outside the neighborhood of the move.
 Definition 4.
@@
 Definition 6.
-Let $X$ be a birack and $R$ a commutative unital ring. A birack bracket is defined by two maps $A,B:X\times X\to R^{\times}$ denoted $A(x,y)=A_{x,y}$ and $B(x,y)=B_{x,y}$ such that
+Let $X$ be a birack and $R$ a commutative unital ring.
+A birack bracket is defined by two maps $A,B:X\times X\to R^{\times}$ from ordered pairs of elements of $X$ to units in $R$ denoted $A(x,y)=A_{x,y}$ and $B(x,y)=B_{x,y}$ such that
   • (i)
@@
 This definition is motivated by the observation that if $\sigma:X\to X$ is a birack endomorphism and $v_{1}$ is a valid birack coloring of a diagram of $L$, then applying $\sigma$ to each color in $v_{1}$ yields another valid coloring, $v_{2}=v_{1}\sigma$.
-As in previous cases such as [3, 5, 7] etc., we observe that since the ingredients for $\mathcal{Q}_{X,S}(L)$ are all invariant under framed Reidemeister moves, so is $\mathcal{Q}_{X,s}(L)$.
+As in previous cases such as [3, 5, 7] etc., we observe that since the ingredients for $\mathcal{BRQ}_{X,S}(L)$ are all invariant under framed Reidemeister moves, so is $\mathcal{BRQ}_{X,s}(L)$.
 That is, we have:
 Theorem 3.
-For any finite biquandle $X$ and $S\subseteq\mathrm{Hom}(X,X)$, the quiver $\mathcal{Q}_{X,S}(L)$ is an invariant of framed isotopy.
+For any finite birack $X$ and $S\subseteq\mathrm{Hom}(X,X)$, the quiver $\mathcal{Q}_{X,S}(L)$ is an invariant of framed isotopy.
 Next, following earlier work such as [7], we come to our primary new definition.
@@
 Then the birack bracket quiver $\mathcal{BRQ}_{X,S}^{\beta}(K)$ has vertex set $V$ and edge set $E$.
-Let $P$ the set of maximal non-repeating paths in the quiver.
+Let $P$ the set of maximal non-repeating paths (i.e., ordered sequences of edges with each edge terminating at the initial vertex of the next edge such that no edge is included more than once and such that the path is not a sub-path of any other path) in the quiver.
 We then decategorify this birack bracket quiver $\mathcal{BRQ}_{X,S}^{\beta}(K)$ to obtain the birack bracket in-degree polynomial